Aerospace

A signal travels a distance of 75 ft in the time it takes to complete 1 cycle. What is its frequency?

1 m 5 3.28 ft

a) 13Mhz

b) 13.12Mhz

c) 14Mhz

d) Cannot be determined
Answer #1

Answer: b

Explanation: 1m=3.28ft

75/3.28=22.86m

Frequency= 300/22.86 =13.12Mhz.


answered by: alibaser